# Hybrid quantum-classical systems: statistics, entropy, microcanonical ensemble and its connection to the canonical ensemble

### v1

@@ III Description of the hybrid microcanonical ensemble
The equilibrium condition for these ensembles can be imposed in terms of the entropy. According to the principles of thermodynamics, an equilibrium distribution will maximize the entropy of the system. Thus, a maximum entropy computation can be performed to determine the equilibrium distributions..
In this section, we review entropy definitions presented in [13], in order to apply the maximum entropy principle to determine the canonical and microcanonical ensembles of hybrid systems.
@@ III.1 Entropy definitions
$S_{N}[\rho]=-\mathrm{Tr}(\rho\log(\rho)),\quad\rho\in\mathcal{P}_{Q}.$ (22)
@@
In the case of hybrid systems, the appropriate entropy function that presents the desirable extensivity properties for these ensembles is [13]:
@@
Notice that, within the bundle picture of figure 1, the hybrid entropy computes for each $\xi\in M_{C}$ the entropy of the quantum subsystem by von Neumann’s formula (22); if the result is interpreted as the classical entropy for a system with classical state $\xi$, then the extensivity property motivates the integration over all possible values of $\xi$, resulting in (23).
The hybrid density (23) reduces to (21) or (22) for those ensembles for which one of the subsystems is pure.
III.2 The hybrid microcanonical ensemble from a maximum entropy principle
@@
In order to apply the maximum entropy principle, let us consider the hybrid Hamiltonian observable $\hat{H}\in\hat{\mathcal{O}}_{H}$, determining the energy of the system. As described above, the hybrid Hamiltonian defines, at each $\xi\in M_{C}$, a quantum observable $\hat{H}(\xi)\in\mathcal{O}_{Q}$ for which the eigenvalue equation can be stated. Assuming a discrete finite spectrum of the Hamiltonian at every $\xi$, the eigenvalue equation is
@@
where $N_{i}^{\xi}$ denotes the multiplicity of the $i$-th eigenvalue at $\xi$. The $\xi$-dependence of both eigenvalues and eigenvectors is made explicit, as both can, independently, vary along with $\xi$. The resulting eigenvectors define the adiabatic basis $b_{\hat{H}}(\xi)=\{|u_{i,j}^{\xi}\rangle\}$, a basis of $\mathcal{H}$ that is different for every classical state $\xi$.
@@
$\mathcal{B}_{\hat{H}}=\left\{(\xi,\pi_{i,j}^{\xi})\in M_{H}\mid\pi_{i,j}^{\xi}=|u_{i,j}^{\xi}\rangle\langle u_{i,j}^{\xi}|,|u_{i,j}^{\xi}\rangle\in b_{\hat{H}}(\xi)\right\}$ (25)
By (16), hybrid densities can be fully described by determining the contribution of each element in $\mathcal{B}_{\hat{H}}$. In order to describe the hybrid microcanonical ensemble, only those elements in $\mathcal{B}_{\hat{H}}$ with energy in a certain interval $E,E+\epsilon]$ should be allowed, which results in the following subset:
$\mathcal{A}_{\hat{H}}^{E,\epsilon}=\left\{\left(\xi,\pi_{i,j}^{\xi}\right)\in\mathcal{B}_{\hat{H}}\mid\mathrm{Tr}\left(\hat{H}(\xi)\pi_{i,j}^{\xi}\right)\in[E,E+\epsilon]\right\}$ (26)
Let $\iota(\xi)$ denote the (possibly empty) set of pairs $(i,j)$ such that $(\xi,\pi_{i,j}^{\xi})\in\mathcal{A}_{\hat{H}}^{E,\epsilon}$. Physically, this set indicates which elements of the adiabatic basis at $\xi$ have energy within the allowed interval. Restricting (16) to (26) gives the following expression for the hybrid density matrix describing the hybrid microcanonical ensemble at every $\xi\in M_{C}$:
$\hat{\rho}_{\mu}(\xi)=\left\{\begin{aligned} &\sum_{(i,j)\in\iota(\xi)}p_{i,j}^{\xi}\pi_{i,j}^{\xi},&&\iota(\xi)\neq\emptyset\\ &0,&&\iota(\xi)=\emptyset\end{aligned}\right$. (27)
The coefficients $p_{i,j}^{\xi}$ represent the probabilities associated to states $(\xi,\pi_{i,j}^{\xi})$, and have to be determined by a maximum entropy principle. In order to be globally defined, we will consider that these coefficients exist for every $\xi\in M_{C}$, being equal to zero if the the corresponding hybrid state does not have energy within the allowed interval, i.e. $p_{i,j}^{\xi}=0$ if either $\iota(\xi)=\emptyset$ or $(i,j)\notin\iota(\xi)$. With these values, the hybrid density matrix (27) can be simply written as
@@
For the sake of simplicity, we will assume in the following that these coefficients can be defined by a generalized function on $M_{H}$. In such a case, the classical marginal distribution can be obtained by taking the trace of (28) to define
@@
Evaluating the functional $\widetilde{S}_{H}$ on the proposal for the microcanonical ensemble(27), and as $\mathrm{Tr}(\pi)=1$ for every rank-1 projector $\pi\in M_{Q}$, the expression of $\widetilde{S}_{H}$ depends only on the probabilities $P_{(\xi,\pi)}$, with integration only over $\xi$ values with contribution to $\hat{\rho}_{\mu}$.
@@
A maximum of $\widetilde{S}_{H}$ is achieved by requiring its variation $\delta\widetilde{S}_{H}$ to be zero, which depends on the variations $\delta p_{i,j}^{\xi}$ of the probabilities $p_{i,j}^{\xi}$, defined for all $\xi\in M_{C}$
@@
$\hat{\rho}_{\mu}(\xi)=\frac{1}{\Omega}\left\{\begin{aligned} &\sum_{(i,j)\in\iota(\xi)}\pi_{i,j}^{\xi},&&\iota(\xi)\neq\emptyset\\ &0,&&\iota(\xi)=\emptyset\end{aligned}\right$. (35)
with $\Omega$ a normalization constant such that $\int_{M_{C}}\mathrm{Tr}(\hat{\rho}_{\mu}(\xi))d\mu_{C}=1$.
@@
    $\begin{gathered}\mathcal{B}_{\hat{H}}=\{(\xi,1)\in M_{H}\}=M_{H},\\ \mathcal{A}_{\hat{H}}^{E,\epsilon}=\{(\xi,1)\in M_{H}\mid h(\xi)\in[E,E+\epsilon]\}\end{gathered}$ (36)
@@
    $\hat{\rho}_{\mu}(\xi)=\left\{\begin{aligned} &\frac{1}{\Omega},&&h(\xi)\in[E,E+\epsilon]\\ &0,&&\mbox{otherwise}\end{aligned}\right$. (37)
@@
    $\mathcal{B}_{\hat{H}}=\{\xi_{0}\}\times\{\pi_{i,j}=|u_{i,j}\rangle\langle u_{i,j}|\mid|u_{i,j}\rangle\in b_{H}\}$ (38)
@@
Lastly, the proposed formalism can be easily extended to a continuous spectrum of energy. It is enough to consider in (27), instead of a sum, an integral over the relevant subset of $M_{Q}$ for the desired energy interval, as it is usually done in the purely quantum case [18].
@@ IV The microcanonical ensemble in the limit of small range of energies
In classical statistical mechanics, it is the case that the microcanonical ensemble, given for an energy interval $E,E+\epsilon]$ by (37), can be represented in the $\epsilon\to 0$ limit by a Dirac delta distribution $\delta\left(H_{C}(\xi)-E\right)$, where $H_{C}\in\mathcal{O}_{C}$ is the Hamiltonian function of the classical system. This limit is usually understood as a sequence where $\epsilon$ can be taken arbitrarily small, and the limit $\epsilon\rightarrow 0$ is well defined in the space $\mathcal{P}_{C}$ of probability distributions. The physical interpretation of this limit is that, if the resolution on the measurement of the energy of an isolated system increases arbitrarily, the microcanonical ensemble is always qualitatively the same, just with narrower energy localization.
In quantum mechanics this is not always the case. In particular, given the quantum microcanonical ensemble (39) for an energy interval $E,E+\epsilon]$ of a quantum system with Hamiltonian $H_{Q}\in\mathcal{O}_{Q}$, if $\epsilon<e_{i}-E$ for every $e_{i}\in{\rm spec}(H_{Q})$, then the quantum microcanonical ensemble is empty, assigning null probability to all events. In fact, if the energy spectrum is discrete, for small enough $\epsilon>0$ the quantum microcanonical ensemble is empty for almost all $E\in\mathbb{R}$; if $\epsilon$ is large enough to accommodate a single energy state, then the microcanonical ensemble only assigns non-zero probability to its eigenstates, while for large values of $\epsilon$ the microcanonical ensemble is an equiprobable mixture of eigenstates of the Hamiltonian whose eigenvalues lay within the energy range. This shows that the quantum microcanonical ensemble cannot be understood in general as a convergent sequence with similar qualitative features to that of classical statistics. Notice, though, that this analysis makes sense, only, for isolated (and hence idealized) systems.
Hybrid statistics, while sharing many relevant characteristics with quantum statistics, nonetheless shows the same qualitative behavior as classical statistics in the limit of small energy values of the microcanonical ensemble in most relevant examples. In order to show this property, let us consider a simple hybrid system in which its classical variables $\xi=(R,P)$ represent classical positions $R$ and momenta $P$. Consider also that the hybrid Hamiltonian $\hat{H}\in\hat{O}_{H}$ contains purely classical terms incorporating a kinetic term $\frac{1}{2}P^{2}$ and a potential term $V(R)$, while the description of the quantum subsystem and its interaction with the classical part is contained in an additional term $H_{Q}(R)$ depending parametrically only on $R$:
@@
with $\mathbb{I}$ the identity operator. Notice that the kinetic energy is unbounded from above this is a very common property of hybrid systems in applications and does not affect the states in the adiabatic basis. For every $(R,P)\in M_{C}$, consider the spectrum ${\rm spec}(\hat{H}(R,P))$ of the hybrid Hamiltonian, its ground state $e_{0}(R,P)$ and its absolute minimum value $E_{0}=\min_{(R,P)\in M_{C}}e_{0}(R,P)$. Then, for any $E\geq E_{0}$ and any $\epsilon>0$, consider the hybrid microcanonical ensemble in an interval $E,E+\epsilon]$ defined by (35), which can also be written as
@@
It is clear that the set of eigenstates of the adiabatic basis with energy within this interval is always non-empty regardless of $\epsilon$, as the kinetic term does not modify the basis but increases its energy by any arbitrary amount. If the $\epsilon\to 0$ limit is taken in (41),
@@
with the usual definition of Dirac delta distribution as the limit of an indicator function [44]. Thus, the hybrid microcanonical ensemble for this system has the desirable property of having non trivial statistical behavior for all $\epsilon>0$ and having the limit $\epsilon\to 0$ defined in a distributional sense. Notice that, as degeneracy is taken into account, there exists for each $\xi\in M_{C}$ at most a single state with energy $E$. Thus, for every $\xi\in M_{C}$, either $\hat{\rho}_{\mu}^{0}(\xi)=0$ or it is proportional to a single Dirac delta distribution. On the other hand, as long as only a discrete and finite subset of $M_{C}$ describe hybrid states with energy $E$, the partition function $Z_{0}$ is simply
@@
More generally, instead of (40), consider a hybrid Hamiltonian whose spectrum varies continuously with the classical variables. For $\epsilon>0$, consider the hybrid microcanonical ensemble (35) for an energy interval $E,E+\epsilon]$. If there exists at least one state of the adiabatic basis with energy in this interval, and as long as the corresponding energy does not present a relative extreme within the interval, then due to the continuity of the spectrum the parameter $\epsilon$ can be taken arbitrarily small, and in the $\epsilon\to 0$ limit the hybrid canonical ensemble will be different from zero.
@@
IV.1 The limit classical marginal distribution
Let us now consider the classical marginal distribution in the $\epsilon\to 0$ limit. In this shape, it is most evident that the microcanonical ensemble is the linear combination of tensor products of several classical (in terms of Dirac’s delta distributions) and quantum (with $\epsilon$ small enough to include just one eigenstate in the case of discrete spectrum) microcanonical ensembles. Let us show this by computing in the first the marginal classical distribution of (43)

### v2

@@ III Description of the hybrid microcanonical ensemble
The equilibrium condition for these ensembles can be imposed in terms of the entropy. According to the principles of thermodynamics, an equilibrium distribution will maximize the entropy of the system. Thus, a maximum entropy computation can be performed to determine the equilibrium distributions, as we showed in Ref. [14] for the hybrid canonical ensemble.
In this section, we review entropy definitions presented in Ref. [14], to apply the maximum entropy principle to determine the microcanonical ensemble of hybrid systems.
@@ III.1 Entropy definitions
$S_{N}[\rho]=-\mathrm{Tr}(\rho\log(\rho)),\quad\rho\in\mathcal{P}_{Q},$ (22)
@@
In the case of hybrid systems, the entropy function with extensivity properties for these ensembles is [14]
@@
Notice that, within the bundle picture of Fig. 1, the hybrid entropy computes for each $\xi\in M_{C}$ the entropy of the quantum subsystem by von Neumann’s formula (22); if the result is interpreted as the classical entropy for a system with classical state $\xi$, then the extensivity property motivates the integration over all possible values of $\xi$, resulting in (23).
The hybrid density (23) reduces to Eq. (21) or Eq. (22) for those ensembles for which one of the subsystems is pure.
III.2 Hybrid microcanonical ensemble from a maximum entropy principle
@@
To apply the maximum entropy principle, let us consider the hybrid Hamiltonian observable $\hat{H}\in\hat{\mathcal{O}}_{H}$, determining the energy of the system. As described above, the hybrid Hamiltonian defines, at each $\xi\in M_{C}$, a quantum observable $\hat{H}(\xi)\in\mathcal{O}_{Q}$ for which the eigenvalue equation can be stated. Assuming a discrete finite spectrum of the Hamiltonian at every $\xi$, the eigenvalue equation is
@@
where $N_{i}^{\xi}$ denotes the multiplicity of the $i$th eigenvalue at $\xi$. The $\xi$-dependence of both eigenvalues and eigenvectors is made explicit, as both can, independently, vary along with $\xi$. The resulting eigenvectors define the adiabatic basis $b_{\hat{H}}(\xi)=\{|u_{i,j}^{\xi}\rangle\}$, a basis of $\mathcal{H}$ that is different for every classical state $\xi$.
@@
$\mathcal{B}_{\hat{H}}=\left\{(\xi,\pi_{i,j}^{\xi})\in M_{H}\mid\pi_{i,j}^{\xi}=|u_{i,j}^{\xi}\rangle\langle u_{i,j}^{\xi}|,|u_{i,j}^{\xi}\rangle\in b_{\hat{H}}(\xi)\right\}.$ (25)
By Eq. (16), hybrid densities can be fully described by determining the contribution of each element in $\mathcal{B}_{\hat{H}}$. To describe the hybrid microcanonical ensemble, only those elements in $\mathcal{B}_{\hat{H}}$ with energy in a certain interval $E,E+\epsilon]$ should be allowed, which results in the following subset:
$\mathcal{A}_{\hat{H}}^{E,\epsilon}=\left\{\left(\xi,\pi_{i,j}^{\xi}\right)\in\mathcal{B}_{\hat{H}}\mid\mathrm{Tr}\left(\hat{H}(\xi)\pi_{i,j}^{\xi}\right)\in[E,E+\epsilon]\right\}.$ (26)
Let $\iota(\xi)$ denote the (possibly empty) set of pairs $(i,j)$ such that $(\xi,\pi_{i,j}^{\xi})\in\mathcal{A}_{\hat{H}}^{E,\epsilon}$. Physically, this set indicates which elements of the adiabatic basis at $\xi$ have energy within the allowed interval. Restricting Eq. (16) to Eq. (26) gives the following expression for the hybrid density matrix describing the hybrid microcanonical ensemble at every $\xi\in M_{C}$:
$\hat{\rho}_{\mu}(\xi)=\left\{\begin{aligned} &\sum_{(i,j)\in\iota(\xi)}p_{i,j}^{\xi}\pi_{i,j}^{\xi},&&\iota(\xi)\neq\emptyset,\\ &0,&&\iota(\xi)=\emptyset.\end{aligned}\right$. (27)
The coefficients $p_{i,j}^{\xi}$ represent the probabilities associated to states $(\xi,\pi_{i,j}^{\xi})$, and have to be determined by a maximum entropy principle. To be globally defined, we will consider that these coefficients exist for every $\xi\in M_{C}$, being equal to zero if the the corresponding hybrid state does not have energy within the allowed interval, i.e. $p_{i,j}^{\xi}=0$ if either $\iota(\xi)=\emptyset$ or $(i,j)\notin\iota(\xi)$. With these values, the hybrid density matrix (27) can be simply written as
@@
For the sake of simplicity, we will assume in the following that these coefficients can be defined by a generalized function on $M_{H}$. In such a case, the classical marginal distribution can be obtained by taking the trace of Eq. (28) to define
@@
Evaluating the functional $\widetilde{S}_{H}$ on the proposal for the microcanonical ensemble(27), and as $\mathrm{Tr}(\pi)=1$ for every rank-1 projector $\pi\in M_{Q}$, the expression of $\widetilde{S}_{H}$ depends only on the probabilities $P_{(\xi,\pi)}$, with integration only over $\xi$ values with contribution to $\hat{\rho}_{\mu}$:
@@
A maximum of $\widetilde{S}_{H}$ is achieved by requiring its variation $\delta\widetilde{S}_{H}$ to be zero, which depends on the variations $\delta p_{i,j}^{\xi}$ of the probabilities $p_{i,j}^{\xi}$, defined for all $\xi\in M_{C}$ as
@@
$\hat{\rho}_{\mu}(\xi)=\frac{1}{\Omega}\left\{\begin{aligned} &\sum_{(i,j)\in\iota(\xi)}\pi_{i,j}^{\xi},&&\iota(\xi)\neq\emptyset,\\ &0,&&\iota(\xi)=\emptyset.\end{aligned}\right$. (35)
with $\Omega$ a normalization factor such that $\int_{M_{C}}\mathrm{Tr}(\hat{\rho}_{\mu}(\xi))d\mu_{C}=1$. This factor can be computed as:
$\Omega=\int_{M_{C}}d\mu_{C}|\iota(\xi)|\,$, (36)
where $|\iota(\xi)|$ is the cardinal or number of elements in the set $\iota(\xi)$. Notice that, disregarding the constant required to fix the dimensionality, $\Omega$ corresponds to the microcanonical partition function and allows us to obtain magnitudes such as the entropy as
$S_{H}(\hat{\rho}_{\mu})=\log\left(\Omega\right)\,.$ (37)
@@
    $\begin{gathered}\mathcal{B}_{\hat{H}}=\{(\xi,1)\in M_{H}\}=M_{H},\\ \mathcal{A}_{\hat{H}}^{E,\epsilon}=\{(\xi,1)\in M_{H}\mid h(\xi)\in[E,E+\epsilon]\},\end{gathered}$ (38)
@@
    $\hat{\rho}_{\mu}(\xi)=\left\{\begin{aligned} &\frac{1}{\Omega},&&h(\xi)\in[E,E+\epsilon],\\ &0,&&\mbox{otherwise},\end{aligned}\right$. (39)
@@
    $\mathcal{B}_{\hat{H}}=\{\xi_{0}\}\times\{\pi_{i,j}=|u_{i,j}\rangle\langle u_{i,j}|\mid|u_{i,j}\rangle\in b_{H}\}.$ (40)
@@
Lastly, the proposed formalism can be easily extended to a continuous spectrum of energy. It is enough to consider in Eq. (27), instead of a sum, an integral over the relevant subset of $M_{Q}$ for the desired energy interval, as it is usually done in the purely quantum case [19].
@@ IV The microcanonical ensemble in the limit of small range of energies
In classical statistical mechanics, it is the case that the microcanonical ensemble, given for an energy interval $E,E+\epsilon]$ by Eq. (39), can be represented in the $\epsilon\to 0$ limit by a Dirac $\delta$ distribution $\delta\left(H_{C}(\xi)-E\right)$, where $H_{C}\in\mathcal{O}_{C}$ is the Hamiltonian function of the classical system. This limit is usually understood as a sequence where $\epsilon$ can be taken arbitrarily small, and the limit $\epsilon\rightarrow 0$ is well defined in the space $\mathcal{P}_{C}$ of probability distributions. The physical interpretation of this limit is that, if the resolution on the measurement of the energy of an isolated system increases arbitrarily, the microcanonical ensemble is always qualitatively the same, just with narrower energy localization.
In quantum mechanics this is not always the case. In particular, given the quantum microcanonical ensemble (41) for an energy interval $E,E+\epsilon]$ of a quantum system with Hamiltonian $H_{Q}\in\mathcal{O}_{Q}$, if $\epsilon<e_{i}-E$ for every $e_{i}\in{\rm spec}(H_{Q})$, then the quantum microcanonical ensemble is empty, assigning null probability to all events. In fact, if the energy spectrum is discrete, for small enough $\epsilon>0$ the quantum microcanonical ensemble is empty for almost all $E\in\mathbb{R}$; if $\epsilon$ is large enough to accommodate a single energy state, then the microcanonical ensemble only assigns nonzero probability to its eigenstates, while for large values of $\epsilon$ the microcanonical ensemble is an equiprobable mixture of eigenstates of the Hamiltonian whose eigenvalues lay within the energy range. This shows that the quantum microcanonical ensemble cannot be understood in general as a convergent sequence in the space of density states with similar qualitative features to that of classical statistics. Notice, though, that this analysis makes sense, only, for isolated (and hence idealized) systems. In such a case, assuming that it is possible to perform experiments with arbitrary precision and measuring suitable magnitudes as the variance of the energy, we should expect a discontinuous behavior when the value of $\epsilon$ selects several eigenspaces, just one, or none at all. Had we performed a similar sequence of experiments on a classical system, the variance of the energy would have changed continuously.
Hybrid statistics, while sharing many relevant characteristics with quantum statistics, nonetheless shows the same qualitative behavior as classical statistics in the limit of small energy values of the microcanonical ensemble in most relevant examples. To show this property, let us consider a simple hybrid system in which its classical variables $\xi=(R,P)$ represent classical positions $R$ and momenta $P$. Consider also that the hybrid Hamiltonian $\hat{H}\in\hat{O}_{H}$ contains purely classical terms incorporating a kinetic term $\frac{1}{2}P^{2}$ and a potential term $V(R)$, while the description of the quantum subsystem and its interaction with the classical part is contained in an additional term $H_{Q}(R)$ depending parametrically only on $R$:
@@
with $\mathbb{I}$ the identity operator. Notice that the kinetic energy is unbounded from above this is a very common property of hybrid systems in applications and does not affect the states in the adiabatic basis. For every $(R,P)\in M_{C}$, consider the spectrum ${\rm spec}(\hat{H}(R,P))$ of the hybrid Hamiltonian, its ground state $e_{0}(R,P)$ and its absolute minimum value $E_{0}=\min_{(R,P)\in M_{C}}e_{0}(R,P)$. Then, for any $E\geq E_{0}$ and any $\epsilon>0$, consider the hybrid microcanonical ensemble in an interval $E,E+\epsilon]$ defined by Eq. (35), which can also be written as
@@
It is clear that the set of eigenstates of the adiabatic basis with energy within this interval is always nonempty regardless of $\epsilon$, as the kinetic term does not modify the basis but increases its energy by any arbitrary amount. If the $\epsilon\to 0$ limit is taken in Eq. (43), then
@@
with the usual definition of Dirac $\delta$ distribution as the limit of an indicator function [49].
While the recovery of the $\delta$ function limit in the hybrid case is an interesting theoretical result, one may argue that this property is just a formal idealization introduced in textbooks in the context of classical mechanics. However, we consider that this result has realistic physical implication not only in the value of the limit itself (which relates to the variance of the energy of the ensemble, and is a desirable property for a thermodynamical ensemble), but in the sequencial nature of the limit. The fact that the accuracy of measurements can be arbitrarily improved without qualitative changes is a requirement for the statistical theory to be consistent in its range of application. After all, being able to have compatible results in this sense under increasingly finer measurements is at the core of the appropriate definition of a microcanonical ensemble. In other words, we here prove that, even if different ensembles could be compatible with varying accuracy in the measurements, only the microcanonical ensemble is compatible with any possible accuracy.
In summary, the hybrid microcanonical ensemble for this system has the desirable property of having non trivial statistical behavior for all $\epsilon>0$ and having the limit $\epsilon\to 0$ defined in a distributional sense. Notice that, as degeneracy is taken into account, there exists for each $\xi\in M_{C}$ at most a single state with energy $E$. Thus, for every $\xi\in M_{C}$, either $\hat{\rho}_{\mu}^{0}(\xi)=0$ or it is proportional to a single Dirac $\delta$ distribution. However, as long as only a discrete and finite subset of $M_{C}$ describe hybrid states with energy $E$, the partition function $Z_{0}$ is simply
@@
More generally, instead of Eq. (42), consider a hybrid Hamiltonian whose spectrum varies continuously with the classical variables. For $\epsilon>0$, consider the hybrid microcanonical ensemble (35) for an energy interval $E,E+\epsilon]$. If there exists at least one state of the adiabatic basis with energy in this interval, and as long as the corresponding energy does not present a relative extreme within the interval, then due to the continuity of the spectrum the parameter $\epsilon$ can be taken arbitrarily small, and in the $\epsilon\to 0$ limit the hybrid canonical ensemble will be different from zero. Again, assuming that experiments can be made with arbitrary precision, the variance of the total hybrid energy should exhibit a continious behavior when $\epsilon$ is reduced (for $\epsilon$ small enough so that only one eigenstate contributes at each $\xi$ at most).
@@
IV.1 Limit classical marginal distribution
Let us now consider the classical marginal distribution in the $\epsilon\to 0$ limit. In this shape, it is most evident that the microcanonical ensemble is the linear combination of tensor products of several classical (in terms of Dirac’s $\delta$ distributions) and quantum (with $\epsilon$ small enough to include just one eigenstate in the case of discrete spectrum) microcanonical ensembles. Let us show this by computing in the first the marginal classical distribution of (45)